OpenStudy (anonymous):

if the sum of a number and seven is doubled, the result is two less than the number. find the number.

OpenStudy (anonymous):

2(x + 7) = x - 2 2x + 14 = x - 2 2x - x = -2 - 14 x = -16 check.. 2(-16 + 7) = -16 - 2 2(- 9) = - 18 -18 = -18 (correct) x = -16

OpenStudy (phi):

Do you think you can change the words into the "math" ? sum of a number and seven when you see a vague term like "a number", give it a name (people like to use "x" ) sum of x and 7 now write that as x+7 to continue the sum of a number and seven is doubled (x+7) is doubled doubled means we have twice as much. if we had x we would have 2x in this case, we have (x+7), so now we have 2(x+7)

OpenStudy (phi):

the result is two less than the number "the result is" turns into = (often the "IS" is a clue to put in an equal sign) 2(x+7) = two less than the number change two to 2 and "the number" to x 2(x+7) = 2 less than x you should learn that 2 less than x means x-2 2 more than x would be x+2 so we now have 2(x+7) = x-2 now it is algebra.

OpenStudy (anonymous):

-16

OpenStudy (phi):

Yes, kellie showed the algebra up above.